9) A mail-order company receives an average of five orders per 500 solicitations. If it sends9) out 100 advertisements, find the probability of receiving at least two orders.
Answer #1

9)

rate = 5 /500

\lambda = 5/500 * 100 = 1

X follow poisson distribution

P(X >= 2)

= 1 - P(X <= 1)

= =1 - POISSON(1,1,1)

=0.2642